### v2

# TimeFormer: Capturing Temporal Relationships of Deformable 3D Gaussians for Robust Reconstruction

Abstract
Dynamic scene reconstruction is a long-term challenge in 3D vision. Recent methods extend 3D Gaussian Splatting to dynamic scenes via additional deformation fields and apply explicit constraints like motion flow to guide the deformation. However, they learn motion changes from individual timestamps independently, making it challenging to reconstruct complex scenes, particularly when dealing with violent movement, extreme-shaped geometries, or reflective surfaces. To address the above issue, we design a plug-and-play module called TimeFormer to enable existing deformable 3D Gaussians reconstruction methods with the ability to implicitly model motion patterns from a learning perspective. Specifically, TimeFormer includes a Cross-Temporal Transformer Encoder, which adaptively learns the temporal relationships of deformable 3D Gaussians. Furthermore, we propose a two-stream optimization strategy that transfers the motion knowledge learned from TimeFormer to the base stream during the training phase. This allows us to remove TimeFormer during inference, thereby preserving the original rendering speed. Extensive experiments in the multi-view and monocular dynamic scenes validate qualitative and quantitative improvement brought by TimeFormer. Project Page: https://patrickddj.github.io/TimeFormer/
footnotetext: $\dagger$ Corresponding Author
1 Introduction
High-quality reconstruction of dynamic scenes is significantly challenging in computer vision and graphics, yet has a wide range of potential applications in movie production, virtual reality, and augmented reality. The difficulty stems from factors like occlusions, translucent materials, specular surfaces, and changing topology, all of which are prevalent in dynamic scenes.
Inspired by the success of neural radiance field (NeRF) on static scenes [36, 3, 37], extending NeRF to dynamic scenes [24, 51, 72, 39, 38, 30, 26, 15] has been explored. However, these NeRF-based methods are limited by computationally intensive volume rendering [7], which makes real-time rendering almost impossible. Recently, 3D Gaussian Splatting (3DGS) [22] represents the scene with anisotropic 3D Gaussians and develops a rasterization-based rendering algorithm, which allows real-time rendering through directly projecting 3D Gaussians onto the image plane. Following the revolutionized 3DGS, dynamic Gaussian Splatting methods [27, 9, 68, 49] have been introduced for reconstructing dynamic scenes. These methods typically construct a canonical 3DGS and utilize a deformation field to deform it based on individual timestamps [67, 27, 61].
However, temporal relationships of 3D Gaussians have been poorly investigated. Previous studies have primarily modeled motion patterns using various types of the deformation field, for example, MLPs [34, 2, 33, 44, 67, 19, 71, 54], spatial-temporal planes [61, 32, 10], polynomial functions [27], Fourier series [21], and combinations of these methods [28]. These methods learn motion patterns from independent time input in a vanilla way, neglecting internal cross-time relationships. Some studies introduce the motion flow regularization [23] to explicitly learn motion patterns from neighboring frames [34, 32]. Although these methods promote similar motion between adjacent timestamps, they adopt a local perspective on time series during optimization. This constraint makes it challenging to reconstruct scenes with more complex motion patterns, such as sudden appearances, violent movements, or reflective surfaces.
In this paper, we introduce TimeFormer, a transformer module designed to implicitly learn motion patterns across multiple timestamps. By modeling temporal relationships within a time batch, we aim to provide a global view of the entire time series, enabling the deformation field backbones themselves to capture motion patterns from a learning perspective, as shown in Fig. LABEL:fig:_teaser. Specifically, TimeFormer utilizes a Cross-Temporal Encoder to capture the implicit motion patterns of 3D Gaussians across multiple sampled timestamps using a self-attention mechanism. Moreover, to avoid additional computational costs of TimeFormer during inference, we present a two-stream optimization strategy. By sharing the weights of two deformation fields, we transfer the motion pattern learning from TimeFormer stream to the base stream during training. Therefore, we can eliminate TimeFormer during inference and maintain the original rendering speed. Notably, TimeFormer does not require any prior information and extracts motion patterns solely from RGB supervision, which can be seamlessly adapted to previous deformable 3D Gaussian methods in a plug-and-play manner. Furthermore, TimeFormer promotes faster gradient descent and guides a more efficient canonical space, ultimately increasing FPS during inference, as demonstrated in Fig. 1.
To sum up, our main contributions are as follows:
  • We propose TimeFormer, a transformer module that enhances current deformable 3D Gaussians reconstruction methods in a plug-and-play manner from an automatic learning perspective.
  • The two-stream optimization strategy allows the exclusion of TimeFormer during inference while maintaining and even improving rendering speed.
  • Extensive experiments on real-world datasets validate the effectiveness of TimeFormer, achieving state-of-the-art rendering quality.
Figure 1: TimeFormer guides towards more efficiently distributed canonical space, showing higher FPS and better quality. The results are from “cut lemon” in the HyperNeRF Dataset [39].
2 Related Works
2.1 Dynamic Scene Reconstruction
Dynamic Scene Reconstruction has been extensively researched over many years, with a wealth of studies [20, 74, 39, 67, 66] contributing to current progress. The pioneering neural radiance field [36, 3, 69, 48, 12] has demonstrated photorealistic rendering for novel view synthesis from calibrated multiview images, inspiring extensive approaches [8, 63, 52, 42, 38, 11, 39] to extend NeRF to 4D space-time field for dynamic scene reconstruction. There are majorly two lines of NeRF-based methods for dynamic scene reconstruction: 1) deformation-based methods [52, 42, 38, 11, 39] model deformation changes by using a deformation field to map the queried positions in different timestamps to a canonical space; 2) compact 4D space-time fields [8, 63, 25, 5] take the timestamps, positions, and directions as input to predict color and density. However, NeRF-based approaches are limited to training and rendering speed. Thus, a large number of methods [24, 29, 56, 58, 43, 13, 4, 64, 62, 57, 47, 1, 31] have been proposed to accelerate NeRF-based methods.
Recent 3DGS [22] achieves real-time rendering and dynamic 3DGS approaches [68, 67, 27] emerge. In details, deformation-based methods [34, 2, 33, 44, 67, 19, 71, 54, 61, 32, 10, 27, 21, 28] employ a deformation field to predict per-Gaussian offsets and deform a canonical 3DGS according to different timestamps. Meanwhile, 4DGS [68] and 4D-Rotor [9] add the time dimension to 3D Gaussian, forming a 4D Gaussian representation. TimeFormer focuses on modeling the temporal relationship and is plug-and-play to deformation-based methods [67] and 4D space-time representations [68, 27]. There are also a few attempts [49, 44, 28] in leveraging the nearby motion or flow for dynamic Gaussian reconstruction. For example, 3DGStream [49] proposes a per-frame optimization strategy, which predicts current Gaussian attributes based on previous Gaussian attributes. Meanwhile, Gaussian-Flow [28] presents explicitly model time-dependent residual of each attribute. To model a long video sequence, SWinGS [44] splits a video sequence into multiple sliding windows based on optical flow and uses a deformation-based method to model each sliding window.
2.2 Motion Modeling
Motion prediction in dynamic scene reconstruction from multi-view videos or monocular videos is an ill-posed problem. Early dynamic NeRF [15, 52, 38, 39] directly learns motion patterns from individual timestamps , and several later works propose motion flow regularization terms [15, 26, 46, 51, 30, 40, 55, 45, 70, 35] to promote the learning of cross-time motion patterns. These methods typically utilize 2D prior information (e.g., optical flow) from pre-trained networks [50] to supervise the scene flows within neighboring frames. PREF [46] uses motion predictor to infer current motion based on the previous four frames and propose a self-supervision strategy without optical flow prior. KFD-NeRF [70] models motion patterns by Kalman Filter based on the previous two frames. CT-NeRF [35] employs a cross-attention mechanism from transformer [53] to learn the correlation of conservative frames.
Dynamic 3D Gaussian Splatting (3DGS) methods [17, 59, 16, 73] also utilize optical flow supervision to enhance reconstruction performance. Techniques such as MD-Splatting [10], D3DG [34], and ST-4DGS [23] focus on minimizing both forward and backward Gaussian flow to promote temporal smoothness in Gaussian motions. DN-4DGS [32] incorporates information from previous, current, and next timestamps into the deformation field to capture cross-time motion patterns. However, these methods primarily establish temporal correlations with neighboring timestamps , which limits their ability to address long-term motion changes. Additionally, they introduce extra computational costs during inference, which can reduce rendering speed. In contrast, TimeFormer captures the temporal relationships of Gaussians from a global perspective across the entire time series. Notably, TimeFormer is employed only during the training phase, without any additional computational costs during inference.
3 Preliminary: Deformable 3D Gaussians
3D Gaussians [22] represent the scene with a set of 3D Gaussians, each of which has unique opacity $o\in[0,1]$, center position $\mu\in\mathbb{R}^{3\times 1}$, and covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$. For a position $x\in\mathbb{R}^{3\times 1}$ in 3D space, the corresponding contribution of a 3D Gaussian on it can be formulated as:
$G(x)=o\cdot e^{-\frac{1}{2}(x-\mu)^{\top}\Sigma^{-1}(x-\mu)}.$ (1)
The covariance matrix $\Sigma$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$: $\Sigma=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{S}=\text{diag}([s_{x},s_{y},s_{z}])$ and $\mathbf{R}$ can be transformed from a quaternion $r_{w},r_{x},r_{y},r_{z}]$. Then the 3D Gaussians can be splatted to a 2D camera plane through differential Gaussian splatting. To model the appearance of 3D Gaussians, spherical harmonics (SH) are introduced to define the color $c$. Finally, for each pixel, the rendering results of 3DGS can be derived by calculating the color contribution of all the related Gaussians. This process is known as $\alpha$-blending:
$C=\sum_{i}^{N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
where $c_{i}$, $\alpha_{i}$ represent the color and density computed from the $i$-th 3D Gaussian.
Recent methods [67, 61, 27] utilize deformation fields that extend 3DGS to 4D space, inspired from NeRF-based methods such as D-NeRF [42]. The structure of the deformation field $\mathcal{D}$ can vary among MLP [67], K-Plane [61] and Polynominal [27], while the deformation process can be summarized as follows:
$(\Delta\mu,\Delta r,\Delta s)=\mathcal{D}(\mu,t),$ (3)
where timestamp $t\in\mathcal{T}$, $\mathcal{T}\in\mathbb{R}^{T\times 1}$ contains $T$ linear time inputs, $\mu,r,s$ are the center position, rotation quaternion, and scaling factors of 3D Gaussians, and $\Delta\mu,\Delta r,\Delta s$ are their residuals, respectively. The inputs and outputs of Eq. 3 can vary among different methods, while they share the same framework.
Figure 2: The Framework of Deformable 3D Gaussians Reconstruction with TimeFormer. Existing deformable 3D Gaussians framework usually includes the canonical space and the deformation field (first row), we incorporate TimeFormer to capture cross-time relationships and explore motion patterns implicitly (second row). We share weights of two deformation fields to transfer the learned motion knowledge. This allows us to exclude this Auxiliary Training Module during inference.
Figure 3: The Structure of Cross-Temporal Encoder. We concatenate randomly sampled timestamps to position $\mu$, treating them as special tokens in a sequence. This module is designed to model multi-temporal relationships and produce distinct time-variant position offsets $\Delta p_{0},\dots,\Delta p_{B-1}$.
Figure 4: Data Flow Changes in the Deformation Field. Dashed lines represent new data flow among time samples $t_{0},\dots,t_{B-1}$.
4 Method
In this section, we first provide an overview of the enhanced deformable 3D Gaussians reconstruction with the proposed TimeFormer (Sec. 4.1). We then introduce the TimeFormer in detail (Sec. 4.2) which consists of a Cross-Temporal Encoder and shared deformation fields (Sec. 4.3). We also discuss the insights behind TimeFormer from the view of gradient flow.
4.1 Overview
Previous methods model motion patterns by explicitly learning temporal relationships on individual or neighboring timestamps, failing on those complex scenes containing violent movement or dynamic reflective surfaces. In contrast, we present TimeFormer to enable the deformable 3D Gaussian backbones themselves to model cross-time relationships from an implicit learning perspective. The main framework with the proposed TimeFormer is shown in Fig. 2. Our approach retains standard reconstruction modules, which include (1) 3D Gaussians in the canonical space and (2) a deformation field that applies time-variant transformation. Additionally, TimeFormer is introduced before the deformation field to extract implicit cross-time motion features for each Gaussian through a self-attention mechanism along the time dimension. Moreover, we share the weights of two deformation fields to transfer the motion knowledge from TimeFormer to mitigate the gap between the original branch and TimeFormer branch, which supports real-time rendering without TimeFormer during inference.
4.2 TimeFormer
Cross-Temporal Encoder BatchFormer [18] demonstrates that the attention mechanism helps learn sample relationships from batch dimension, rather than channel and spatial dimentions [53, 6], inspiring us that different timestamps can also be considered as a special time batch. Let $\mathcal{T}_{s}\subset\mathcal{T},\mathcal{T}_{s}=\{t_{i}\}_{i=0}^{B-1}$ denotes randomly sampled timestamps, and let $\mathcal{G}\in\mathbb{R}^{N\times(3+C)}$ denotes Gaussians in the canonical space, where $B$ is the size of time batch, $N$ is the number of Gaussians, $3+C$ means each Gaussian has 3 position channels and $C$ additional channels. As in Fig. 3, all Gaussians’s postions are made into $B$ copies, expanded into $\mathcal{G}_{c}\in\mathbb{R}^{B\times N\times 3}$, and sampled timestamps are made into $N$ copies, expanded into $\mathcal{T^{\prime}}\in\mathbb{R}^{B\times N\times 1}$. Then, we composite $\mathcal{G}_{c}$ and $\mathcal{T^{\prime}}$ together and apply position encoding function $\gamma$ to extract high frequency information, as in Eq. 4:
$\gamma(p)=$ $\left(\sin\left(2^{0}\pi p\right),\cos\left(2^{0}\pi p\right),\right$$\cdots$ (4)
$\left.\sin\left(2^{L-1}\pi p\right),\cos\left(2^{L-1}\pi p\right)\right)$
In our experiments, we set $L=6$ for both positions $x$ and time $t$. We treat $\gamma(\mathcal{G}_{c},\gamma(\mathcal{T^{\prime}})]\in\mathbb{R}^{B\times N\times(3\times 2L+2L)}$ as the original input $F_{0}$ to TimeFormer. It can be considered as $N$ sequences of the length $B$, containing $8L$ feature channels. With $M$ transformer encoder layers, for $m^{th}$ layer, intermediate features are calculated through multi-head self-attention(MSA) and MLP blocks. In the final stage, we use a tiny MLP to transform the last encoded features $F_{M-1}$ into offset $\mathcal{O}\in\mathbb{R}^{B\times N\times 3}$ in the linear space:
$\mathcal{O}=MLP(F_{M-1}),\quad\mathcal{G}_{t}=\mathcal{G}_{c}+\mathcal{O}$ (5)
We consider the output from the Cross-Temporal Encoder as a fixing residual term to the original positions to encourage a gradual, steady learning process on motion patterns. Implicit cross-time relationship learning in TimeFormer enables the automatic aggregation of Gaussians with similar variations during optimization, promoting more efficient spatial distribution and accelerating rendering speed.
Gradient Analysis Let $\mathcal{D}$ and $\mathcal{P}$ be the deformation field and TimeFormer respectively, the output of the deformation field $\Delta\mu_{i}$ and partial derivative for $\mu$ in time $t_{i}\in\mathcal{T}_{s}$ are formulated as follows:
$\Delta\mu_{i}=\mathcal{D}(\mu,t_{i}),\quad\frac{\partial\Delta\mu_{i}}{\partial\mu}=\frac{\partial\mathcal{D}(\mu,t_{i})}{\partial\mu}$ (6)
With TimeFormer applied on time batch $\mathcal{T}_{s}$, we reformulate as Eq. 7 and data flow changes as in Fig. 4.
$\Delta\mu_{i}=\mathcal{D}(\mu+\mathcal{P}(\mu,\mathcal{T}_{s}),t_{i})$ (7)
To calculate the partial derivative for $\mu$, we apply the chain rule. Let $a=\mu+\mathcal{P}(\mu,\mathcal{T}_{s})$, then:
$\frac{\partial\Delta\mu_{i}}{\partial\mu}=\frac{\partial\mathcal{D}(a,t_{i})}{\partial a}\cdot\frac{\partial a}{\partial\mu},\quad\frac{\partial a}{\partial\mu}=1+\frac{\partial\mathcal{P}(\mu,\mathcal{T}_{s})}{\partial\mu}$ (8)
$\frac{\partial\Delta\mu_{i}}{\partial\mu}=\frac{\partial\mathcal{D}(a,t_{i})}{\partial a}\cdot\left(1+\frac{\partial\mathcal{P}(\mu,\mathcal{T}_{s})}{\partial\mu}\right)$ (9)
Eq. 6 has a weaker dynamic nature, as it only considers the current timestamp $t_{i}$. In contrast, Eq. 9 incorporates an additional gradient term, $\frac{\partial\mathcal{P}(\mu,\mathcal{T}_{s})}{\partial\mu}$, enabling the current state to be influenced by any past or future states. In other words, $\Delta\mu_{i}$ also optimizes the models according to other timestamps $t_{j}\in\mathcal{T}_{s}(j\neq i)$, which is a significant difference compared to the backward process without TimeFormer. Such design accounts for cross-time attention and allows the models to capture more challenging motion patterns from a global view of the entire time series.
4.3 Shared Deformation Fields
Accounting for additional computation costs of TimeFormer, which can significantly decrease rendering speed, we force the original deformation field and the auxiliary deformation field to share weights for knowledge transferring. Let $\mathcal{V}$ be the camera viewpoints, we apply the shared deformation fields to predict deformed space from both Gaussians in the canonical space and Gaussians from TimeFormer. Then, we apply the splatting algorithm to these two groups of deformed space. We calculate the losses between rendered images and ground truth $\mathcal{I}_{gt}$ as follows:
$\mathcal{L}_{c}=\|Splatting(\mathcal{D}(\mathcal{G}_{c},\mathcal{T}),\mathcal{V})-\mathcal{I}_{gt}\|_{1}$ (10)
$\mathcal{L}_{t}=\|Splatting(\mathcal{D}(\mathcal{G}_{t},\mathcal{T}),\mathcal{V})-\mathcal{I}_{gt}\|_{1}$ (11)
$\mathcal{L}=\lambda_{c}\mathcal{L}_{c}+\lambda_{t}\mathcal{L}_{t}$ (12)
, where $\mathcal{L}_{c},\mathcal{L}_{t}$ represent losses of original branch and TimeFormer branch with $\lambda_{c}>\lambda_{t}$. We use a relatively smaller $\lambda_{t}$ because we find it easy to overfit on the second branch with TimeFormer, causing a degradation in inference quality.
Method Sear Steak Flame Salmon Cut Beef Flame Steak Cook Spinach Coffee Martini Mean
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
K-Plane [14] 32.52 0.971 30.44 0.942 31.82 0.965 32.38 0.970 30.60 0.968 29.99 0.943 31.63 0.960
MixVovels [56] 31.21 0.971 29.92 0.945 31.30 0.965 31.43 0.970 31.61 0.965 29.36 0.946 30.80 0.960
GS4D [68] 32.92 0.953 26.39 0.897 33.08 0.959 33.81 0.967 32.77 0.956 25.23 0.884 30.07 0.936
4D-Rotor [9] 32.86 0.956 28.25 0.913 33.14 0.952 31.61 0.953 32.56 0.949 27.95 0.908 31.06 0.938
4DGS [61] 32.49 0.949 28.92 0.917 32.90 0.956 32.51 0.954 32.46 0.948 27.34 0.903 31.10 0.938
+TimeFormer 33.38 0.955 29.33 0.924 33.11 0.957 33.25 0.953 33.03 0.949 28.93 0.910 31.84 0.941
STGS [27] 33.71 0.962 28.21 0.921 33.52 0.958 33.46 0.963 33.13 0.955 27.71 0.915 31.62 0.946
+TimeFormer 34.34 0.965 29.13 0.924 33.57 0.958 34.04 0.964 33.45 0.956 28.83 0.917 32.23 0.947
Table 1: Quantitative Comparisons on N3DV Dataset [25]. We use bold font to indicate the improvement, statistics of K-Plane [14] and MixVoxels [56] are from their original paper, while we calculate metrics of all Gaussian-based methods by running their official codes.
Figure 5: Visualization of Comparisons on N3DV Dataset [25].
Figure 6: Per-Frame PSNR on N3DV Dataset [25]. TimeFormer improves the reconstruction quality, especially at the beginning and end time series.
Figure 7: Comparions of Convergence Speed. We use the same batch size for 4DGS [61] and calculate the loss of both branches in our method.
Method NN $\downarrow$ Training $\downarrow$ FPS $\uparrow$
STGS [27] 172.1 k 41 min 82.1
+TimeFormer 148.9 k 78 min 88.9
4DGS [61] 145.9 k 52 min 31.7
+TimeFormer 113.1 k 94 min 37.9
(a) Quantitative comparisons on N3DV Dataset [25].
Method NN $\downarrow$ Training $\downarrow$ FPS $\uparrow$
DeformGS [67] 169.6 k 25 min 30.1
+TimeFormer 82.9 k 35 min 58.9
4DGS [61] 172.3 k 32 min 35.7
+TimeFormer 135.5 k 48 min 40.9
(b) Quantitative comparisons on HyperNeRF Dataset [39].
Table 2: Comparisons of Gaussian Number, Training Time & FPS.
Method Broom Chicken Cut Lemon Torchco Peel Banana Hand Mean
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DeformGS [67] 20.74 0.322 26.32 0.786 27.94 0.714 27.4 0.877 26.38 0.836 27.79 0.78 26.09 0.719
+TimeFormer 21.01 0.324 26.55 0.792 30.01 0.780 27.55 0.883 27.24 0.852 29.79 0.848 27.03 0.747
4DGS[61] 21.53 0.351 26.82 0.797 29.72 0.763 27.45 0.883 27.82 0.844 29.52 0.841 27.14 0.747
+TimeFormer 22.84 0.401 27.06 0.801 30.23 0.778 29.48 0.902 28.21 0.853 30.38 0.862 28.03 0.766
Table 3: Quantitative Comparisons on HyperNeRF Dataset [39].
Figure 8: Visualization of Comparisons on HyperNeRF Dataset [39].
Method Press Plate Basin Sieve Bell Cup Mean
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
HyperNeRF [39] 25.4 0.873 18.1 0.714 20.2 0.829 25.0 0.909 24.0 0.884 24.1 0.896 22.8 0.851
NeRF-DS [65] 26.4 0.911 20.8 0.867 20.3 0.868 26.1 0.935 23.3 0.872 24.5 0.916 23.57 0.895
DeformGS[67] 25.68 0.866 20.82 0.812 19.87 0.804 25.71 0.881 24.92 0.854 24.52 0.897 23.59 0.852
+TimeFormer 26.29 0.867 20.90 0.817 19.93 0.805 26.26 0.891 25.90 0.873 25.16 0.903 24.10 0.859
Table 4: Quantitative Comparisons on NeRF-DS Dataset [65].
Figure 9: Visualization of Comparisons on NeRF-DS Dataset [65]. Compared with original results from DeformGS [67], TimeFormer presents a clearer visual effect on dynamic objects with specular surfaces.
Figure 10: Motion Visualization on HyperNeRF Dataset [39].
5 Experiment
5.1 Implementation Details
Our implementation is tested on a single A100 GPU. We use $M=4$ transformer encoder layers in TimeFormer and the number of time samples is $B=4$. We set $\lambda_{c}=1$ and $\lambda_{t}=0.8$ to prevent overfitting on TimeFormer. We assess our experimental results using image quality metrics, including peak-signal-to-noise ratio (PSNR) and structural similarity index (SSIM [60]).
Baselines & Datasets. We apply TimeFormer to 4DGS [68], STGS [27] on multi-view videos, e.g., N3DV Dataset [25], and evaluate TimeFormer on 4DGS [68] and DeformGS [67] on monocular dynamic videos, e.g., HyperNeRF Dataset [39], respectively. We also utilize NeRF-DS Dataset [65] to demonstrate the robustness of TimeFormer on moving objects with specular surfaces.
5.2 Experimental Comparisons
Multi-View Video Dataset. In Tab. 1, TimeFormer improves the reconstruction quality (e.g., PSNR) of original 4DGS [61] and STGS [27] by 0.74 and 0.61, outperforming all baselines methods. We extract the frames at an initial time for scene Cook Spanish in Fig. 5, and TimeFormer can reconstruct a clearer geometry and specular effects compared to the original results.
To have a more comprehensive overview of TimeFormer’s capacity on different temporal stages, we collect PSNR for 300 frames of all six scenes from N3DV Dataset [25] and calculate the average per-frame PSNR. As in Fig. 6, the reconstruction quality increases more significantly at the beginning and end, and such balanced reconstruction results are attributed to TimeFormer’s cross-temporal attention mechanism on the whole time series.
Monocular Video Dataset. Tab. 3 shows the enhancements achieved by TimeFormer on the HyperNeRF Dataset. It’s observed that TimeFormer increases the PSNR by 0.89 for 4DGS [61] and 0.94 for DeformGS [67], and the SSIM by 0.019 and 0.028, respectively. Fig. 8 additionally provides visualization of the improvement in the image quality: on the left, TimeFormer generates a clearer outline between fingers, while on the right the texture on the front of the broom is clearer. TimeFormer also eliminates artifacts and reveals a clear structure of the broom.
Dynamic Reflective Video Dataset. Fig. 9 additionally demonstrates that TimeFormer can work well on scenes with dynamic specular objects on NeRF-DS Dataset [65]. TimeFormer can reconstruct a cleaner appearance on reflective material (e.g., a metallic cup, a glass bottle), and this improvement is attributed to the attention mechanism which enables the deformation field to detect changes on specular surfaces from a learning perspective automatically.
Analysis of FPS & Canonical Space. We also find that TimeFormer achieves more efficient representation with much fewer points in the final canonical space, thereby achieving a higher FPS in Tab. 2. We attribute the phenomenon to the reason that TimeFormer guides faster gradient descent and promotes the spatial layout of Gaussian points during optimization. Interestingly, in Fig. 7, the L1 loss of both branches in our method decreases faster and shows almost the same convergence trend during training, exceeding the optimization speed of the original method. This eliminates many redundant points and guides the canonical space toward a more efficient distribution, as in Fig. 1. Moreover, Fig. 7 also proves that the cross-time relationship learned in TimeFormer has been successfully transferred to the base branch, as introduced in Sec. 4.3.
[…]
In Fig. 10a, TimeFormer identifies the blade’s motion as a rigid deformation, showing a sharper outline and consistent motion, especially as it contacts the initially static lemon. This demonstrates TimeFormer ’s ability to distinguish between moving and static objects. In Fig. 10b, TimeFormer captures subtle motion on the spray gun and clear flame flickering on the beef. Fig. 10c shows that TimeFormer also detects the violent movement of the elongated broom, where 4DGS [61] fails, proving its robustness in handling extremely geometries.
5.3 Ablation Studies
Cross-Temporal Encoder. Tab. 5 demonstrates that the performance of TimeFormer is not sensitive to time batch $B$ and transformer encoder layer $M$. In Tab. 5(b), we explore two sampling methods: random sampling and continuous sampling of timestamps. Both strategies demonstrate improvements in PSNR and other metrics compared to the baseline. Notably, the random sampling method yields more significant enhancements. We hypothesize that this phenomenon occurs because the random sampling method allows for more effective propagation of dynamic behavior modeling across the entire time series.
Shared Deformation Fields. In Tab. 5(b), the image quality during inference suffers a considerable decrease if we train two deformation fields without sharing their weights. Fig. 7 also demonstrates the effectiveness of the two-stream strategy and TimeFormer’s ability of re-balancing.
Setting Random Sampling Continuous Sampling
[…]
BB=2 25.89 0.875 0.117 25.73 0.871 0.124
BB=4 26.06 0.871 0.119 25.78 0.874 0.117
BB=6 25.99 0.87 0.123 25.73 0.865 0.125
(a) Ablation Studies on Batch Size and Sampling Strategies.
[…]
Baseline 25.44 0.867 0.125
M=1 25.96 0.874 0.117
M=2 25.90 0.874 0.120
M=3 26.05 0.876 0.112
M=4 26.04 0.873 0.119
w/o Shared 24.81 0.852 0.132
(b) Ablation Results. $M$ is the number of encoder layers, “w/o Shared” means not using shared weights in two deformation fields.
Table 5: Ablation Results. The results are from three scenes press, sieve and bell on NeRF-DS Dataset [65].
6 Conclusion
We propose TimeFormer, a Transformer module that is plug-and-play to existing deformable 3D Gaussians methods and enhances reconstruction results without additional computational budget. TimeFormer enables deformation fields to implicitly model complex motion patterns from a learning perspective. In addition, we design a two-stream optimization strategy to transfer the learned motion knowledge from TimeFormer to the original deformation branch. This allows us to remove TimeFormer during inference and thus maintain the same inference speed as the original methods. Extensive experiments demonstrate TimeFormer effectively facilitates the reconstruction of three state-of-the-art deformable 3D Gaussians Splatting methods among three datasets, and illustrate the improvement on reconstructing complex scenes containing violent movement, extreme-shaped geometries, or reflective surfaces.
Limitations TimeFormer may produce overly smooth textures for objects with intricate details.
Appendix A Implementation
In this section, we provide the Pytorch [41] code of TimeFormer in Alg. 1 and two-stream optimization strategy in Alg. 2, to clarify the framework in Fig. 2.
Alg. 1 is an additional explanation for Fig. 3, including three parts: 1) Position Encoding (PE), 2) Definition of Transformer encoder, 3) Definition of tiny MLP. Note that we use a shared TimeFormer on all Gaussians in the canonical space.
The Transformer Encoder receives input structured as [seq_len, seq_batch, channel], and we input x structured as [$B$, NN, 4], where $B$ is the size of time batch, $N$ is the number of Gaussians and 4 means 3 position channel and 1 channel for the timestamp. Alg. 2 shows how we construct input to time.
class TimeFormer ($n$.module):
# d_in: here is 4, (x, y, z, t)
# L: frequency of Position Encoding (PE) $\gamma$
def __init__(self, d_in, L, nhead, d_hidden, n_layer):
[…]
self.PE, PE_ch = get_PE(L=L, d_in=d_in)
# define Cross-Temporal Encoder
layer = $n$.TransformerEncoderLayer(PE_ch, nhead, d_hidden, activation=$n$.functional.tanh)
self.encoder = $n$.TransformerEncoder(layer, n_layer)
# define Tiny MLP
self.mlp = $n$.Linear(PE_ch, 3)
# x: [seq_len, seq_batch, channel]
def forward(self, x):
PE_x = self.PE(x)
h = self.encoder(PE_x)
h = self.mlp(h)
return h
Algorithm 1 Implementation of TimeFormer.
# timeformer: TimeFormer
# deform: Shared Deformation Field
# N: number of Gaussians
# GS: Gaussians in the canonical space
# B: size of time batch
B = 4
lambda_t = 0.8
# Vs, Ts: sampled cameras and timestamps
# images_gt: sampled GT images
Vs, Ts, images_gt = random.sample(Dataset, B)
# construct input to TimeFormer
# [N, 3] => [B, N, 3]
G_expanded = GS.xyz.unsqueeze(0).expand(B, -1, -1)
# [B] => [B, N, 1]
T_expanded = Ts.unsqueeze(1).expand(-1, N).unsqueeze(2)
# src: [B(seq_len), N(seq_batch), 4(channel)]
src = torch.cat((G_expanded, T_expanded), dim=2)
# offset_t: [B, N, 3]
offset_t = timeformer(src)
# use iterations to save cuda memory
loss = 0.0
for i in range(B):
# original branch
# simplified: (xyz, t) => d_xyz
d_xyz = deform(GS.xyz, Ts[i])
image = splatting(GS, Vs[i], d_xyz=d_xyz)
loss += L1(image, images_gt[i])
# TimeFormer branch: use offset_t[i,:,:]
d_xyz_t = deform(GS.xyz+offset_t[i,:,:], Ts[i])
image_t = splatting(GS, Vs[i], d_xyz=d_xyz_t)
loss += lambda_t * L1(image_t, images_gt[i])
# Optimize shared deformation field, TimeFormer and Gaussians in the canonical space
loss.backward()
deform.optimizer.step()
timeformer.optimizer.step()
GS.optimizer.step()
Algorithm 2 Two-Stream Optimization Strategy.
In Alg. 2, we introduce the process in a time batch in detail. We first construct input to TimeFormer by concatenating Gaussian positions and sampled time stamps, as in Sec. 4.2. Besides the original branch where the deformation function is directly performed on the canonical space, we add another TimeFormer branch. TimeFormer calculates prior offsets “offset_t” via cross-time relationships before the deformation field. These two branches are optimized at the same time during training, while the TimeFormer branch can be removed during inference.
Appendix B Analysis on Canonical Space & FPS
Apart from Fig. 1, we provide more results in Fig. 11 and Fig. 12, as a further illustration on TimeFormer’s capability to reduce Gaussians in the canonical space and improve inference speed. TimeFormer promotes more efficient spatial distribution of Gaussians in the canonical space, leading to improvements in reconstruction quality while simultaneously eliminating a substantial number of redundant Gaussians compared to baseline methods.
Figure 11: Comparisons of Canonical Space, FPS on Hypernerf Dataset [39].
Figure 12: Comparisons of Canonical Space, FPS on NeRF-DS Dataset [65].
Setting Bell Press Sieve Mean
[…]
Baseline 24.92 0.854 0.126 25.68 0.866 0.141 25.71 0.881 0.108 25.44 0.867 0.125
M=1 25.67 0.872 0.097 26.17 0.867 0.141 26.04 0.884 0.113 25.96 0.874 0.117
M=2 25.46 0.872 0.104 26.01 0.866 0.139 26.22 0.883 0.116 25.90 0.874 0.120
M=3 25.87 0.875 0.095 26.29 0.865 0.138 26.00 0.887 0.104 26.05 0.876 0.112
M=4 25.71 0.870 0.103 26.09 0.865 0.14 26.33 0.885 0.115 26.04 0.873 0.119
w/o Shared 24.12 0.832 0.137 25.01 0.854 0.146 25.29 0.869 0.113 24.81 0.852 0.132
Table 6: Ablation Results of three scenes press, sieve and bell on NeRF-DS Dataset [65].
Appendix C Slider Window Demo
We provide additional comparison results of baseline methods with TimeFormer. We strongly recommend opening the following website demos in folder website_demos and using the “slider window” to see the improvements brought by TimeFormer more clearly.
  • Overview: index.html
  • 4DGS+TimeFormer on HyperNeRF Dataset: index_4DGS_hypernerf.html
  • DeformGS+TimeFormer on HyperNeRF Dataset: index_deformGS_hypernerf.html
  • DeformGS+TimeFormer on NeRF-DS Dataset: index_deformGS_nerfds.html
Appendix D Ablation Studies
We provide more detailed ablation results on three scenes on NeRF-DS Dataset [65], as in Tab. 6. This further illustrates that TimeFormer is not sensitive to the number of transformer encoder layers $M$. However, we observe a significant decrease in reconstruction quality on all scenes without shared deformation fields.